TEXFormat

A GUI for selecting TEX format files

Version 2.00

September 14, 1993

by Richard Bödi

Abstract

TEXFormat is a program that enables you to select TEX format files easily. The program scans the directory where your TEX format files reside and creates an array of radiobuttons of the appropriate size. Moreover the program shows the filenotes of the format files making it easier to remember the purposes of the format files. There are two versions of the program. A very flexible one based an Stefan Stuntz’s MagicalUserInterface (MUI) and a less nice non-MUI version (of the same functionality).

Contents

1. Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2. Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3. Running TEXFormat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1. Running TEXFormat from the CLI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2. Running TEXFormat from Workbench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.2.1. CURRENTFORMAT Tooltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2.2. FORMATPATH Tooltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2.3. PUBSCREEN Tooltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2.4. PATTERN Tooltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4. Selecting a TEX format file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

5. Menu description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

6. Filenotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

7. Required software and hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1. Copyright

TEXFormat is a “freely distributable copyrighted software”, i e. everyone may copy and use it, but the copyright remains by myself (Richard Bödi).

TEXFormat may not be sold commercially without permission. It may not be part of any PD series except Fred Fish’s AmigaLibDisks. Permission to modify the software is granted, but not the right to distribute the modified code or only part of the files.

Thanks go to Stefan Stuntz for his MagicalUserInterface. The MUI version of TEXFormat is my first step in programming with MUI and I think I really like it. My advise for every Amiga owner is: get a registrated version of MUI and watch out for new MUI based applications which surely will appear in the near future. Last not least I want to thank Fred Fish for the continuing release of his library disks.

2. Introduction

Have you ever worked with several different TEX format files when typesetting with TEX? One format file for your letters, another one for program documentations, and many others for typesetting the documents of your friends? Do you have to switch between format files containing hyphenation tables for different languages? If your answer to one of these questions is ’Yes’, you will surely admit that setting an environment variable is not the most comfortable way for selecting TEX format files. This is why I wrote TEXFormat. This program enables you to select TEX format files easily via an array of radiobuttons.

3. Running TEXFormat

The TEXFormat program can either be executed in a shell using parameters or from Workbench using ToolTypes.

3.1. Running TEXFormat from the CLI

Format: TeXFormat CURRENTFORMAT=<env-variable>

                  FORMATPATH=<path for TEX formats>
                  [PUBSCREEN=<PubScreen>]
                  [PATTERN=<file pattern for format files>]

Pattern: TeXFormat CURRENTFORMAT/K, FORMATPATH/K, PUBS-CREEN/K, PATTERN/K

Description of the command line options:

CURRENTFORMAT Specifies the environment variable which contains the name of the format files that TEX should use.

              Example for PasTEX: CURRENTFORMAT=texformat.

FORMATPATH The full path to your TEX format files.

              Example : FORMATPATH=tex:formats.

PUBSCREEN This keyword is only available for the non-MUI version. Specifies the name of the public screen on which TEXFormat should open its window. Default is the default public screen. Example: SCREEN=ShowDVI-PubScr opens the TEXFormat window on the ShowDVI previewer screen (ShowDVI is the previewer of the PasTeX implementation). The public screen for the MUI version can be determined by using the screen manager of the MUI preference program, see the documentation of MUI.

PATTERN Specifies the file pattern for TEX format files.

              Example : PATTERN=#?.fmt.

3.2. Running TEXFormat from Workbench

When running TEXFormat from Worbench the parameters are specified by ToolTypes which can be edited by the Workbench’s Info command.

3.2.1. CURRENTFORMAT Tooltype

Specifies the environment variable which contains the name of the format files that TEX should use.

3.2.2. FORMATPATH Tooltype

The full path to your TEX format files.

3.2.3. PUBSCREEN Tooltype

This keyword is only available for the non-MUI version.

Specifies the name of the public screen on which TEXFormat should open its window. Default is the default public screen. The public screen for the MUI version can be determined by using the screen manager of the MUI preference program.

3.2.4. PATTERN Tooltype

Specifies the file pattern for TEX format files.

4. Selecting a TEX format file

To select a TEX format file, simply click on the corresponding radiobutton. The environment variable specified by the CURRENTFORMAT parameter is set immediately after this operation. So you don’t have to quit the program to set the TEX format file. The 'SAVE' button copies the current setting to the environment archive EnvArc:. You can quit the program either by pressing the 'DONE' button, by selecting Quit from the menu, by closing TEXFormat’s window, or by pressing AMIGA-Q (or the key ’D’ in the MUI version).

5. Menu description

The TEXFormat program has only one menu which is the File menu. It has only two items, namely the About and the Quit item. I think the purpose of both of these items is clear enough to omit an explanation.

6. Filenotes

In addition to the selection of format files, TEXFormat shows the filenotes of the format files. Let’s assume we have three format files plain, gplain and lplain. Since we want to be more verbose, we store a comment to each of these files in a filenote, for example

filenote plain "Plain TEX"
filenote gplain "Plain TEX, german hyphenation"
filenote lplain "LaTEX".

After that, when starting the non-MUI version of TEXFormat, we see three radiobuttons labelled with

plain    Plain TEX
gplain   Plain TEX, german hyphenation
lplain   LaTEX

The MUI version of TEXFormat behaves a bit different. It shows three radiobuttons labelled with

plain
gplain
lplain

Only the filenote of the selected format file is shown in a text field below the radiobuttons. But now, the filenotes may contain multiple lines by simply adding a C-style ’\n’. For example, when setting

filenote gplain "Plain TEX\ngerman hyphenation"

we obtain the two lines ’Plain TEX’ and ’german hyphenation’. Note that this feature is only available for the MUI version.

7. Required software and hardware

• An Amiga, of course.

• Kickstart 2.04 or higher.

• A TEX implementation.

• A TEX file to typeset.

• A TEXnician (not required).

If you have any remarks (if you like the program or if not) or suggestions for improvements feel free to contact me:

Richard A. Bödi Mathematisches Institut Universität Tübingen Auf der Morgenstelle 10 72076 Tübingen GERMANY

e-mail (Internet) : mmisa01@mailserv.zdv.uni-tuebingen.de